OpenStudy (anonymous):

i just need to know a method to do this. like quadratic formula but this isnt that x^3+6x^2+50x=20x^2+10x

OpenStudy (kc_kennylau):

1. Put all things to the left hand side 2. Factor out x 3. See what you're left with

OpenStudy (anonymous):

nicely explained kc

OpenStudy (anonymous):

i got x^3-14x-40x=0

OpenStudy (kc_kennylau):

@perl will you stop doing everything

OpenStudy (anonymous):

x(x^2 - 14x + 40) = 0 (x - 4)(x - 10) = 0 x - 4 = 0 x = -4 x - 10 = 0 x = 10

OpenStudy (kc_kennylau):

you too @kelliegirl33

OpenStudy (perl):

x^3+6x^2+50x = 20x^2+10x x^3+6x^2+50x - 20x^2 -10x=0 x^3 -14x^2 + 40x = 0 x( x^2 -14x + 40 ) = 0 x = 0 or x^2 -14x + 40 = 0 x = 0 or (x-10)(x-4)=0 x=0 or (x-10) = 0 or (x-4) = 0 x= 0 or x = 10 or x = 4

OpenStudy (anonymous):

my bad...sorry bout that

OpenStudy (anonymous):

so i got x=0, x=4, or x=10 as my final answer

OpenStudy (kc_kennylau):

@kelliegirl33 it's fine :)

OpenStudy (anonymous):

x = 4 and x = 10...I messed up before and put -4, but it is 4.

OpenStudy (anonymous):

@kc_kennylau ....I will do better next time and try to explain instead of just giving answers

OpenStudy (kc_kennylau):

@kelliegirl33 lolz ok :)